OpenStudy (anonymous):

Determine algebraically whether the function is even, odd, or neither even nor odd. Odd Even Neither

OpenStudy (anonymous):

function is \[f(x)=x+\frac{ 12 }{ x }\]

hartnn (hartnn):

Find f(-x) by replacing x in f(x) by -x. If f(-x) = f(x), then the function is even If f(-x) = -f(x), then the function is odd else neither.

OpenStudy (anonymous):

\[f(-x)=-x-\frac{12}{x}=-(x+\frac{12}{x})=-f(x)\]

OpenStudy (anonymous):

so is an odd function?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

how about \[f(x)=x+x^2\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

neither?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yey!! thank you